# QuaLLM: An LLM-based Framework to Extract Quantitative Insights from Online Forums

### v1

@@ 4.3. Evaluation
We randomly sampled 125 submissions and all the 2,511 associated comments, jointly annotated by two researchers, to create a reference list of concerns. The two researchers then jointly evaluated the LLM outputs against the reference list of concerns, and resolved any differences through discussions. We obtained a factuality score of 0.55 and a completeness score of 0.78. Lower factuality was primarily due to the LLM identifying concerns part of the input context, but not directly related to AI and algorithmic decisions. Moreover, the lower factuality score was less concerning as outputs not factual (i.e., identified by the LLM but not by humans) were mostly filtered out into the “Other” category during the classification step and excluded from any further analysis.
4.3.2. Classification and Prevalence Prompts.
@@
Two researchers jointly annotated 100 LLM-classified concerns to establish ground truth labels, assigning one label per concern. They compared LLM outputs with these labels, achieving a consensus accuracy of 0.74. The achieved accuracy is significant because misclassified samples, excluding those labeled as ”Other,” can still contribute to the analysis. This is because concerns might fit into multiple categories depending on the analytical perspective. Even if placed in a different category, these concerns remain valuable for later aggregation and prevalence prompts. Using a similar methodology, researchers obtained an accuracy of 0.82 for the prevalence prompts’ output.
4.3.3. Aggregation Prompt.
@@
We employed BERTopic for topic modeling (see Appendix D for hyperparameter details). Given the large volume of concerns (4K-25K) per primary theme, human evaluation on a subset would not accurately capture the task’s nuances. Therefore, we adopted this well-established computational method for evaluation. We fit a topic model on the entire text (title+description) associated with each primary theme to obtain the top 5 and 10 topics. Subsequently, we identified the most similar topic for each of the 5 sub-themes. The results demonstrated a distinctness of 0.95, coverage(1) of 0.80, and coverage(2) of 1.00, indicating that the LLMs’ outputs were both distinct and well-represented among the most frequent topics.
5. Discussion of Broader Impacts
5.1. Scaling online forum analysis with LLMs
@@
QuaLLM presents a novel framework that leverages the power of LLMs to analyze vast amounts of unstructured text data efficiently and translate to a form which resembles community survey responses. Beyond analyzing online discussions in rideshare communities on Reddit, this methodology has broad applicability across numerous domains, across any platform, where large amounts of unstructured text is prevalent and traditional positivist quantitative methods (Park et al., 2020) such as topic models through LDA, may fall short. We now discuss a few domains where our approach can aid data analysis:
(1) Healthcare Forums::
@@
(5) Financial Market Sentiments::
@@
5.2. QuaLLM’s evaluation reveals strengths and necessitates further refinement
Due to the unique nature of our pipeline approach, which combines several tasks into a cohesive workflow, unlike the most closely related work by Pham et al. (2023) where LLMs are used for the single task of topic generation, there are no established baselines or benchmarks available for direct end-to-end comparison and evaluation. Topic models might offer a comparison for the generation prompt outputs, yet they require significant human effort to translate outputs to human readable forms, unlike our method where the LLM autonomously generates topics, making baselining challenging.
This absence of precedent opens questions regarding effective baselining and benchmarking for performance evaluation. As a result, human evaluation serves as the gold standard for evaluation, ensuring the validity of our findings.
5.3. LLMs can be used to assist researchers and complement existing methods
Some researchers highlight limitations in LLMs’ understanding of complex human emotions and experiences, pointing to a lack of context-sensitivity (Bender et al., 2021; Alkaissi and McFarlane, 2023; Rudolph et al., 2023). Others (Bano et al., 2023) argue LLMs cannot match the innate human ability to perform certain tasks in inexplicable ways, described in conceptual frameworks for knowledge creation (Li and Gao, 2003; Collins, 2005). This gap raises doubts about LLMs’ capability for interpretation crucial to fields such as anthropology (Malinowski, 1929) and sociology(Weber, 1949), challenging their application within the qualitative interpretivist research paradigm (Gichuru, 2017; Glaser et al., 1968; Denzin, 2001).
On the other hand, advocates for LLM-enhanced research suggest LLMs align with human outputs and in some cases can replace human judgement (Dwivedi et al., 2023; Bano et al., 2023; Chew et al., 2023; Xiao et al., 2023; Dai et al., 2023; Tai et al., 2023; Dunivin, 2024; Pham et al., 2023). Further, they find integrating AI with human expertise can improve the scope and scale of analyses, facilitating efficient processing of vast, unstructured datasets.
Our research stands in the middle, underscores the potential of LLMs to help researchers engage with data in a way that can supply intrinsic meaning at scale, while stressing the need to recognize their limitations and the continued importance of human oversight and interpretation.
Broadly, QuaLLM can also be used in mixed methods study designs to complement data from interviews and focus groups. However, safeguards, including successive filtering and human evaluation, are necessary to ensure validity of LLM outputs, indicating a complementary shift rather than a complete replacement of human interpretation.
6. Limitations
Interpretivist vs Positivist trade-offs::
@@ 6. Limitations
Small scale human evaluation::
We acknowledge the limited generalizability of our human evaluation due to the relatively small sample size and its restriction to specific pipeline stages.
Hyperparameter-sensitive topic modeling::
BERTopic performance is very sensitive to choice of hyperparameters and this can result in significant variation in results based on the chosen hyperparameter values.
Reproducibility concerns::
@@
7. Conclusion
Our study presents a new framework using LLMs consisting of a prompting methodology and evaluation strategy for analysis of online forum data, showcased by analyzing over one million rideshare subreddit comments—the largest analysis to date. This approach highlights LLMs’ capability to efficiently process large unstructured datasets establishing a precedent for future AI-assisted positivist quantitative research. But does an increase in data equate to enhanced understanding? Echoing Agnew et al. (2024)’s sentiments, we believe further research is essential to determine the applicability and utility of AI-assisted research methods.
@@ Acknowledgements
We acknowledge the Azure Cloud Computing Grant from CSML at Princeton which enabled access to Microsoft Azure compute resources. We thank Sayash Kapoor for insightful discussions on the conceptual and normative implications of LLMs and prompting strategies. We are grateful to Vinayshekhar Bannihatti Kumar for his valuable inputs on defining evaluation metrics. We thank Sunnie Kim for her contributions on connections to Explainable AI. We appreciate Elizabeth Anne Watkins for providing valuable critiques from an interpretivist perspective, which helped reframe our work. We thank Yuhan Liu, Beza Desta, and Kiyosu Maeda from the Princeton HCI group for their annotations. We are grateful to Hope Schroeder for their feedback on evaluations and broad discussions on applications. We thank Prof. Arvind Narayanan and Mihir Kshirsagar from Princeton CITP for their insights on the generalization of this work and its policy implications. We thank Prof. Brian Keegan for pointing us to the Reddit data source. Finally, we thank the anonymous CSCW reviewers for their valuable feedback, which helped frame the larger paper of which this is a part.
@@ Appendix A Prompts
GPT-4 Generation Prompt GPT-4 Classification Prompt GPT-4 Aggregation and Ranking Prompt GPT-4 Prevalence Prompt
@@ Appendix B Findings
Table 1. Most representative concerns on Reddit relevant to the need for enhancing transparency and explainability (Total = 24,721).
@@
Table 2. Most representative concerns on Reddit relevant to the need for enhancing predictability and worker agency (Total = 12,728).
@@
Table 3. Most representative concerns on Reddit relevant to the need for better safety and more time (Total = 6,144).
@@
Table 4. Most representative concerns on Reddit relevant to the need for ensuring fairness and non-discrimination (Total = 4,280).
Appendix C Financial Expenditure of Using LLMs
@@
Table 5. GPT4-Turbo Token Usage and Associated Costs
The access to the LLM and associated compute resources through Microsoft Azure was jointly funded through a partnership between our university and Microsoft. The total expenditure for the GPT4-Turbo usage was \$1,662.30. This funding enabled us to leverage the capabilities of GPT4-Turbo to successfully meet our research objectives.
Appendix D BERTopic Hyperparameters
1# Load CSV data
2df = pd.read_csv(filename)
3
4# Concatenate title and description into one text column
5df[’text’] = df[’title’] + " " + df[’description’]
6
7dim_model = UMAP(n_neighbors=15,
8 n_components=10,
9 min_dist=0.0,
10 metric=’cosine’,
11 random_state=42)
12
13# Initialize BERTopic
14topic_model = BERTopic(
15 nr_topics="auto",
16 min_topic_size=100,
17 n_gram_range=(1, 2),
18 umap_model=dim_model
19)
20
21# Fit BERTopic on the concatenated texts from the CSV
22topic_model.fit(df[’text’])
@@ Appendix E IRR Measurements
For the Classification and Prevalence stages, which associated a label to a given text input, we calculated inter-rater reliability (IRR). Three trained annotators annotated 100 randomly chosen samples each for both stages. We calculated human-human IRR and human-LLM IRR using Fleiss’ Kappa. To calculate human-LLM IRR, we generated the human labels by taking the majority vote of the three annotators. If no clear majority existed, i.e., all three annotators chose different labels, we labeled the class as ’Other’. The results were as follows: For the Classification stage, the human-human IRR was 0.59, and the human-LLM IRR was 0.54. For the Prevalence stage, the human-human IRR was 0.63, and the human-LLM IRR was 0.61. These results suggest good agreement according to the interpretation guidelines by Fleiss et al. (2013). Crucially, the human-human and human-LLM IRR values differed by less than 0.05, demonstrating LLMs can substitute human labels.

### v2

@@ 4.3 Evaluation
We randomly sampled 125 submissions and all the 2,511 associated comments, jointly annotated by two researchers, to create a reference list of concerns. The two researchers then jointly evaluated the LLM outputs against the reference list of concerns and resolved any differences through discussions. We obtained a factuality score of 0.55 and a completeness score of 0.78. Lower factuality was primarily due to the LLM identifying concerns that are part of the input context but not directly related to AI and algorithmic decisions. Moreover, the lower factuality score was less concerning as outputs that are not factual (i.e., identified by the LLM but not by humans) were mostly filtered out into the “Other” category during the classification step and excluded from any further analysis.
4.3.2 Classification and Prevalence Prompts.
@@
Two researchers jointly annotated 100 LLM-classified concerns to establish ground truth labels, assigning one label per concern. They compared LLM outputs with these labels, achieving a consensus accuracy of 0.74. The achieved accuracy is significant because misclassified samples, excluding those labeled as “Other,” can still contribute to the analysis. This is because concerns might fit into multiple categories depending on the analytical perspective. These concerns remain valuable for later aggregation and prevalence prompts, even if placed in a different category. Using a similar method, researchers obtained an accuracy of 0.82 for the prevalence prompts’ output.
4.3.3 Aggregation Prompt.
@@
We employed BERTopic for topic modeling (see Appendix C for hyperparameter details). Given the large volume of concerns (4K-25K) per top-level theme, human evaluation on a subset would not accurately capture the task’s nuances. Therefore, we adopted this well-established computational method for evaluation. We fit a topic model on the entire text (title+description) associated with each top-level theme to obtain the top 5 and 10 topics. Subsequently, we identified the most similar topic for each of the 5 sub-themes. The results demonstrated a distinctness of 0.80, coverage(1) of 0.95, and coverage(2) of 1.00, indicating that the LLMs’ outputs were distinct and well-represented among the most frequent topics.
5 Conclusion
Our study introduces a new LLM-based framework with a prompting method and evaluation strategy to analyze online forum data, demonstrated through the largest analysis of over 1 million rideshare subreddit comments to date. The methodology’s recursive nature demonstrates its versatility, e.g., the Aggregation and Prevalence prompts can function as Generation and Classification at the sub-theme level, and initial filtering can reframe as classifying concerns’ relevance to the overall theme. This approach highlights LLMs’ capability to efficiently process large unstructured datasets, establishing a precedent for future AI-assisted positivist quantitative research. But does an increase in data equate to enhanced understanding? Echoing Agnew et al. (2024)’s sentiments, we believe further research is essential.
Ethical and Broader Impacts Statement
Scaling online forum analysis with LLMs
@@ 5 Conclusion
QuaLLM presents a novel framework that leverages the power of LLMs to analyze vast amounts of unstructured text data efficiently and translate it to a form that resembles community survey responses. Beyond analyzing online discussions in rideshare communities on Reddit, this methodology has broad applicability across numerous domains, across any platform, where large amounts of unstructured text are prevalent and traditional positivist quantitative methodologies Park et al. (2020) such as topic models through LDA, may fall short.
Healthcare Forums:
@@
Financial Market Sentiments:
@@
QuaLLM’s evaluation reveals strengths and necessitates further refinement
Due to the unique nature of our pipeline approach, which combines several tasks into a cohesive workflow, there are no established baselines or benchmarks available for direct end-to-end comparison and evaluation. Topic models might offer a comparison for the generation prompt outputs, yet they require significant human effort to translate outputs to human-readable forms, unlike our methodology, where the LLM autonomously generates topics, making baselining challenging.
This absence of precedent opens questions regarding effective baselining and benchmarking for performance evaluation. As a result, human evaluation serves as the gold standard for evaluation.
LLMs can be used to assist researchers and complement existing methods
Some researchers highlight limitations in LLMs’ understanding of complex human emotions and experiences, pointing to a lack of context-sensitivity Bender et al. (2021); Alkaissi and McFarlane (2023); Rudolph et al. (2023). Others Bano et al. (2023) argue LLMs cannot match the innate human ability to perform specific tasks in inexplicable ways, described in conceptual frameworks for knowledge creation Li and Gao (2003); Collins (2005). This gap raises doubts about LLMs’ capability for interpretation crucial to fields such as anthropology Malinowski (1929) and sociology Weber (1949), challenging their application within the qualitative interpretivist research paradigm  Gichuru (2017); Glaser et al. (1968); Denzin (2001).
On the other hand, advocates for LLM-enhanced research suggest LLMs align with human outputs and may replace human judgement Dwivedi et al. (2023); Bano et al. (2023); Chew et al. (2023); Xiao et al. (2023); Dai et al. (2023); Tai et al. (2023); Dunivin (2024). Further, they find integrating AI with human expertise can improve the scope and scale of analyses, facilitating efficient processing of vast, unstructured datasets.
Our research stands in the middle and underscores the potential of LLMs to help researchers engage with data in a way that can supply intrinsic meaning at scale while stressing the need to recognize their limitations and the continued importance of human oversight and interpretation.
Broadly, QuaLLM can also be used in mixed methods study designs to complement data from interviews and focus groups with adequate safeguards and evaluations.
Limitations
Interpretivist vs. Positivist trade-offs:
@@
Small scale human evaluation:
We acknowledge the limited generalizability of our human evaluation due to the relatively small sample size
Hyperparameter-sensitive topic modeling:
BERTopic performance is very sensitive to the choice of hyperparameters, and this can result in significant variation in results based on the chosen hyperparameter values.
Reproducibility concerns:
@@
Bias in LLMs:
LLM outputs are influenced by biases introduced during training, alignment processes, and prompt design, which may affect what concerns are highlighted, favor structured language over informal text, and introduce potential socio-economic or stylistic biases; additionally, API errors (e.g., due to profanity) may create selection bias in the analyzed data.
Misinformation in Online Communities:
Misinformation in online forums can distort analysis, and the spread of false information may reflect user biases or socio-economic factors. However, the noise may smoothen out when aggregated over millions of posts, reducing its overall impact.
@@ Acknowledgements
We acknowledge the Azure Cloud Computing Grant from CSML at Princeton, which enabled access to Microsoft Azure compute resources. We thank Sayash Kapoor for insightful discussions on the conceptual and normative implications of LLMs and prompting strategies. We are grateful to Vinayshekhar Bannihatti Kumar for his valuable input on defining evaluation metrics. We thank Sunnie Kim for her contributions on connections to Explainable AI. We appreciate Elizabeth Anne Watkins for providing valuable critiques from an interpretivist perspective, which helped reframe our work. We thank Yuhan Liu, Beza Desta, and Kiyosu Maeda from the Princeton HCI group for their annotations. We are grateful to Hope Schroeder for her feedback on evaluations and broad discussions on applications. We thank Prof. Arvind Narayanan and Mihir Kshirsagar from Princeton CITP for their insights on the generalization of this work and its policy implications. We thank Prof. Brian Keegan for pointing us to the Reddit data source. Finally, we thank the anonymous CSCW and FAccT reviewers for their valuable feedback, which helped frame the larger paper of which this is a part.
@@ Appendix A Prompts
Listing 1: GPT-4 Generation Prompt
1Analyze a set of JSON objects, each representing a submission from the r/UberDrivers and r/LyftDrivers subreddits. For each JSON object, you will find the following information:
2
3Submission Title: The title of the Reddit post.
4Submission Body: The main content or message of the post.
5Timestamp: The date and time when the post was submitted.
6Group Key: A unique identifier that is common across all 5 submissions in the dataset.
7Comments: A list of all comments made on the submission.
8
9Generate a list of the most frequently occurring and impactful concerns due to AI and algorithmic platform features discussed by drivers within the input context.
10
11Step 1: Identify mentions in submission bodies, titles, and/or comments about concerns that pertain to a lack of knowledge or available information regarding the platforms algorithms for drivers, including but not limited to fares, routes, incentive programs, driver preferences, etc.
12
13Step 2: Group similar concerns across comments and submissions to ensure a mutually exclusive list of concerns and avoid redundancy. For example, multiple mentions of fare calculation issues should be grouped under a single concern.
14
15Step 3: From the grouped concerns, select the most representative quote for each concern. Ensure the quote clearly illustrates the specific concern due to AI and algorithmic platform features.
16
17Step 4: Assess which concerns are mentioned most frequently and have the most significant impact on drivers.
18
19Step 5: Create a list of these concerns in a JSON format. Each entry should include (with these specific field names):
20"title": The title of the concern
21"description": A brief description (10-20 words)
22"quote": The selected representative quote.
23
24Step 6: Ensure the final list is concise, precise, and specifically addresses drivers concerns due to AI and algorithmic platform features present regarding the platforms algorithms and policies. Include only those concerns found in the input context without generalizing based on prior or outside knowledge.
25
26Step 7: Group any similar concerns to avoid redundancy. If there are no concerns, output "No concerns". Do not generate any other text.
Listing 2: GPT-4 Classification Prompt
1Task: Analyze a list of rideshare drivers transparency concerns. Each concern should be evaluated and categorized into one of the following five categories. For each concern listed, assign exactly one letter that corresponds to its most appropriate classification in the same order as the original list (preserving the serial number of the concern).
2
3Categories:
4
5A. Need for Enhancing Transparency and Explainability: Concerns due to the lack of information and explanations about AI and algorithmic features, required for drivers to do their work. These include opaque trip details, unclear surge boundaries, bonus and quest clarity issues, lack of wage breakdown, platform take rate and clarity on how prices and bonuses are calculated, and the impact of ratings on metrics.
6
7B. Need for Greater Predictability and Worker Agency: Concerns due to significant variation in AI and algorithmic features leading to reduced worker agency and diminished predictability of work conditions. These include unpredictable wages, misleading destination filter, immense variation in surge prices and quest matches across drivers and location, and deceptive trip offers.
8
9C. Need for Better Safety and More Time: Concerns focused on safety risks and time pressures due to the AI and algorithmic features. These include dangerous multitasking, compromised route safety, unattainable quests, and acceptance rate concerns, emphasizing the need for improved safety and time management in app design.
10
11D. Need for Ensuring Fairness and Non-Discrimination: Concerns related to algorithmic wage discrimination. These include unequal pay for similar work, earnings below the prevailing minimum wage, influence of the demographic characteristics of drivers and riders on earnings, all caused by opaque AI and algorithmic features.
12
13E. Other Concerns: Any concerns that do not fit into the above categories.
14
15Output Format: Present the analysis in a dictionary format with the serial number of the concern as the key and the classification as the value, preserving the order of the concerns.
16
17{1: A, 2: B, 3: C}
18
19Note: Ensure that each concern is classified under only one of the aforementioned categories and that there is one classification corresponding to each concern in the input, e.g., for 400 transparency concerns, there should be 400 items in the dictionary. Do not generate any other text.
Listing 3: GPT-4 Aggregation and Ranking Prompt
1The data contains a list of concerns of rideshare drivers obtained from discussions on the r/uberdrivers and r/lyftdrivers subreddits. Each concern title and its description is present on consecutive lines.
2
3These concerns are relevant to <insert category description from Classification Prompt>
4
5Identify the 5 most frequently occurring themes of concerns. Ensure each concern is sufficiently different from the others on the list. Dont repeat the same concern in the list. If there are similar concerns, group them and find another one.
6
7Provide the output in a rank ordered format:
8{concern_rank: <a number between 1-5>, concern_title: string, concern_description: string of 10-20 words}
Listing 4: GPT-4 Prevalence Prompt
1Classify each line which contains the title and description of a concern from rideshare drivers on Reddit into the following 6 categories:
2
3<list the 5 categories A to E from the aggregation prompt output>
4F: Other
5
6Output Format: Present the analysis in a dictionary format with the serial number of the concern as the key and the classification as the value, preserving the order of the concerns.
7
8{1: A, 2: B, 3: C}
9
10Note: Ensure that each concern is classified under only one of the aforementioned categories and that there is one classification corresponding to each concern in the input, e.g., for 400 concerns, there should be 400 items in the dictionary. Do not generate any other text.
@@ Appendix B Findings
Table 2: Most representative concerns on Reddit relevant to the need for enhancing transparency and explainability (Total = 24,721).
@@
Table 3: Most representative concerns on Reddit relevant to the need for enhancing predictability and worker agency (Total = 12,728).
@@
Table 4: Most representative concerns on Reddit relevant to the need for better safety and more time (Total = 6,144).
@@
Table 5: Most representative concerns on Reddit relevant to the need for ensuring fairness and non-discrimination (Total = 4,280).
Appendix C BERTopic Hyperparameters
# Load CSV data
df = pd.read_csv(filename)
# Concatenate title and description into one text column
df[’text’] = df[’title’] + " " + df[’description’]
dim_model = UMAP(n_neighbors=15,
n_components=10,
min_dist=0.0,
metric=’cosine’,
random_state=42)
# Initialize BERTopic
topic_model = BERTopic(
nr_topics="auto",
min_topic_size=100,
n_gram_range=(1, 2),
umap_model=dim_model
# Fit BERTopic on the concatenated texts from the CSV
topic_model.fit(df[’text’])
Appendix D Financial Expenditure of Using LLMs
@@
Table 6: GPT4-Turbo Token Usage and Associated Costs
We spent a total of \$1,662.30 for using GPT4-Turbo via Microsoft Azure, which our university helped fund.
@@ Appendix E IRR Measurements
For the Classification and Prevalence stages, which associated a label with a given text input, we calculated inter-rater reliability (IRR). Three trained annotators annotated 100 randomly chosen samples each for both stages. We calculated human-human IRR and human-LLM IRR using Fleiss’ Kappa. To calculate human-LLM IRR, we generated the human labels by taking the majority vote of the three annotators. If no clear majority existed, i.e., all three annotators chose different labels, we labeled the class as ’Other’. The results were as follows: For the Classification stage, the human-human IRR was 0.59, and the human-LLM IRR was 0.54. For the Prevalence stage, the human-human IRR was 0.63, and the human-LLM IRR was 0.61. These results suggest good agreement according to the interpretation guidelines by Fleiss et al. (2013). Crucially, the human-human and human-LLM IRR values differed by less than 0.05, demonstrating LLMs can substitute human labels.